A Bibliography of Publications of Nicholas Ian Gould

Nicholas Ian Gould Rutherford Appleton Laboratory

OX11 0QX England

Tel: 235 445801 FAX: (44) 0235 446626

E-mail: nimg@letterbox.rl.ac.uk (Internet)

17 November 2023 Version 1.37

Abstract

This bibliography records publications of Nicholas Ian Gould.

Title word cross-reference

l[1] [GOT03a]. CUTEst [GOT15]. P^{4} [ADGR90, DAGR87]. P^{5} [ADGR90, DAGR87].

-penalty [GOT03a].

14th [Ame94].

65th [GLT04b].

7th [DMW06].

89h [CGT89a].

90 [GOT03c]. ’94 [Ame94]. 95 [GT03a, GT07].

absolute [GN98]. absolute-value [GN98]. accurate [GM81, Gou86]. Active [Gou03]. Active-Set [Gou03]. Adaptive [CGT11a, CGT11b, CGJR16, CGT12a, CGT12c, GPT12]. al. [DAGR87]. Algebra [DGDG97b, DGDG97a]. Algorithm [CGT91e, DAGR87, BCG^{+}10, BGNW04, CGT12a, CGT92c, CGT92h, CGT92i, CGST94a, CGST96a, CGT97b, CGT97a, CGT97c, CGT00a, CGOT00, FGL^{+}02, FGF13, GGM^{+}82b, Gou91a, GLRT98, GT03c, GLT04a, GT06, GPT12, GOP16]. Algorithms [BGNW05, CGT12d, CGST93b, CGST93a, CGST93c, ADD^{+}94, ADGR90, BGNW03, CG84, CGT88a, CGT89a, CGT92b, CGST93d, CGST96b, CGJR16, GT99, GOST01, GL03, GLT04b, GOST05]. alternative [CGT94a]. Anal [CGT89a]. Analysis [GS10, DW97, GS09]. Announcement [CGT92a]. application [BBG^{+}12, GT03b]. Applications [CGT11c, CGT90b, DGSW10, GLT04b]. Applied [Ame94, DMW06]. approach [CGT94c]. approximate [GS98]. approximate-inverse [GS98]. April [DGDG97b]. Arbitrary [CGT20]. Arbitrary-Order [CGT20]. Arising [GHN01, CGT94a, GM81]. Art [DMW06, DW97]. Atlanta [Ame94]. Augmented [CGT91e, CGST93b, CGST93a, CGST93c, BDG94, CGST96a, CGJR16, DGP94].

B [GOT02]. barrier [CGT92c, CGT94a, CGT97b, CGT97c, GT99]. between [BCG^{+}97a, BCG^{+}97b]. bi [Gou12]. bi-projection [Gou12]. Bibliography [GOT00]. binary [BBG^{+}12]. birthday [GLT04b]. Block [DDG98]. Blocks [GS10, GS09]. bound [CGT00a, FGF13]. Bounds [CGT20, CGT91e, CGT12b, CGT88a, CGT88b, CGT89a, CGT92c, CGT92h, CGT92i, CGT97b, CGT97a, CGT97c, GGM^{+}82b]. branch [FGF13]. buckling [BBG^{+}12].

Calculations [CGT97d, ACD^{+}93]. Case [CGT20, CGT11b]. Chebyshev [CGT13a]. class [CGT88a, CGT88b, CGT89a, CGST93d]. classes [Gou88]. Combination [CGST93a, CGST93c, CGST93b, CGST96a]. comparison [BCG^{+}97a, BCG^{+}97b]. Complete [Gou91b, GS03, GHS05a, BCG^{+}97b, CGT92d]. Complexity [CGT10, CGT11c, CGT12b, CGT12d, CGT14b, CGT15, CGT20, CGT11b, CGT12a, CGT12c, CGT13b, CGT13a, CGT14a]. Componentwise [GOST02]. Composite [CGT11c]. comprehensive [CGT91a, CGT91b]. Computation [Ame94, GLR14]. Computing [ACD^{+}93, DMW06]. conditioned [Gou99a, Gou00]. conditions [Gou85].

Congress [Ame94]. Constrained [BCGT95, CGT14b, CGT15, CGT92f, GHN01, Gou03, ACD^{+}93, BGNW04, CGT13b, CGT14a, CGT93, CGT94d, Gou89, Gou99b, GOT03b, GOT15, BCGT93]. Constraint [KGW00, DGSW05, DGW06, DGSW07]. Constraints [CGT20, CGT91e, CGST93b, CGST93a, CGST93c, CGT97d, BGG15, BBG^{+}12, CGT12a, CGT92b, CGT92c, CGT92h, CGT92i, CGST93d, CGST94a, CGST96a, CGST96b, CGT97b, CGT97a, CGT97c, CGT00a, GGM^{+}82b, GR89]. context [CGT94c]. continuous [FGF13]. Convergence [BCG^{+}10, BGNW05, CGT91c, CGT92b, CGST93b, CGST93a, CGST93c, CGST94a, CGST96a, CGST96b, GR10a, GR10b, GLR15, BGNW03, CGT11a, CGT88a, CGT89a, CGST93d, FGL^{+}02, Gou89, GT99, GOST01, GOST02, GT03c, GT06]. Convergent [CGT91e, CGT92c, CGT92h, CGT92i, CGT97b, CGT97a, CGT97c]. convex [CGT12a, CGT12c, CGT92b, CGST93d, CGST94a, CGST96b, CGOT00, GGM^{+}82a, GGM^{+}84, GT02b, Gou08, GOR13]. core [ADGR90]. Correction [CGT89a]. crash [GR89]. critical [CGT14a]. Cubic [CGT14b, CGT11a, CGT11b, CGT12a, CGT12c, CGT13b, GPT12]. Current [Gou03, CGT94d]. curvature [GLRT00]. CUTE [BCGT95, BCGT93]. CUTEr [GOT03b].

Data [CGT91d, CGT89c, CGT90c]. decomposition [CGT94c]. Deficient [CGT14b, CGT13b]. definite [CG86]. degenerate [GOR13]. Denmark [DMW06]. Derivative [CGT12d, GR10a, GR10b, CGT11b, GR12]. derivative-evaluation [CGT11b]. Derivative-Free [CGT12d]. derivatives [BBG^{+}12]. Descent [CGT10, CG84]. description [CGT91a, CGT91b]. determination [GM81, Gou86]. differentiable [Gou86, Gou88]. dimer [GOP16]. dimer-type [GOP16]. Direct [DGDG97a, DGDG97b, GS03, GS04, GHS05a, GHS05b, GSH07, SHG04]. direct-solution [GS03]. direction [ACD^{+}93]. directions [CG84, Gou86, GLRT00]. dual [CGT00a, CGOT00, GOST01].

earthquake [GM81]. edge [Gou83a, Gou84a, Gou84b]. Efficient [DLG96]. Element [DLG94, DLG96, DLG97, DLG95, DDLG96, DDLG97]. Element-by-Element [DLG94, DLG96, DLG97, DLG95, DDLG96, DDLG97]. Elimination [Gou91b]. Environment [BCGT95, DGLR87, DGLR90, GOT03b, GOT15, BCGT93]. epicenter [GM81]. Equality [CGST93a, CGST93c, GHN01, BGNW04, CGT92h, CGST93b, CGST96a, CGT00a, Gou85]. equations [ADD^{+}94, DLG95, GOST02, GS03, GLT04a, GS04, GHS05a, GHS05b, GSH07]. Erisman [DAGR87]. Erratum [GT12]. estimates [CGL20]. Euclidean [BCG^{+}10]. Evaluation [CGT11c, CGT12c, CGT14b, CGT15, CGT20, CGT11b, CGT12a, CGT13b, GOT02, GS03, GS04, GHS05a, GHS05b, GSH07, SHG04]. exact [CG87]. existence [Gou85]. experience [CGJR16]. Experiments [BDG94, DGP94, CGT92g, CGT96]. Exploiting [GLRT00, CGT94b]. extrapolated [Gou12].

Factorization [DGSW06, ADGR90, DGW06, DGR^{+}90, DGR^{+}91, GN98]. factorizations [Gou99b]. fast [BGG15, BBG^{+}12, GOST02]. feasibility [GT03a, GT03b, GT07, Gou08, Gou12]. Filter [GLR14, GLR15, FGL^{+}02, GT03a, GT03b, GT03c, GLT04a, GST05, GT06, GT07, GT10, GT12]. filter-trust-region [GT03a, GST05, GT07]. FILTRANE [GT03a, GT07]. Finding [CG86, CGT14a]. findings [SHG04]. First [CGT12d, BBG^{+}12, CGT14a, CGT94c]. First-Order [CGT12d, BBG^{+}12, CGT14a]. Fletcher [GLT04b]. following [GOR13]. Foreword [CGT90d]. format [CGT89c, CGT90c, CGT91d]. Fortran [CGT92e, CGT92j, GT03a, GOT03c, GT07]. Free [CGT12d, GR12]. full [GOST02]. full-rank [GOST02]. Function [CGT11c, CGT11b, CGT12a, CGT13a, CG87, CGT94a, CGT00a, Gou89, GT10, GT12]. function- [CGT11b]. function-evaluation [CGT12a]. functions [CGL20, CGT94c, FGF13, Gou86, Gou88].

GALAHAD [GOT03c]. Gaussian [Gou91b]. General [CGT15, CGT91e, CGST93a, CGST93c, CGT92c, CGT92h, CGT92i, CGST93b, CGST96a, CGT97b, CGT97a, CGT97c, FGL^{+}02, GGM^{+}82b, Gou83b, Gou85, GT03c]. generalized [Gou83a, Gou84a]. generated [CGT91c]. Georgia [Ame94]. Global [CGT88a, CGST93d, CGST93b, CGST93a, FGL^{+}02, GT03c, GT06, GR10a, FGF13, CGT89a]. Globally [CGT91e, CGT92c, CGT92h, CGT92i, CGT97b, CGT97a, CGT97c]. good [Gou08, Gou12]. Gram [GGM^{+}84]. Growth [Gou91b].

held [DGDG97b]. Hellerman [DAGR87]. Hessian [FGF13]. honor [GLT04b]. HSL [GS03, GS04]. hybrid [GT03c].

idea [GT03b]. II [CGT11b]. ILAY [DGDG97b]. ill [Gou99a, Gou00]. ill-conditioned [Gou00]. IMACS [Ame94, Ame94]. Implicit [DGSW06, DGW06]. Implicit-Factorization [DGSW06, DGW06]. Improving [CGT94c]. in-core [ADGR90]. Indefinite [GS10, DGR^{+}90, DGR^{+}91, GS09, KGW00]. inequality [CGT92c, CGT92i, CGT97b, CGT97a, CGT97c]. inexact [CGST93d]. Inexpensive [CGT20]. infinite [CG84, CG87]. inner [CGT92h, CGT92i, CGT97a]. Input [CGT91d, CGT89c, CGT90c]. Institute [Ame94]. Intensive [CGT92d]. Interior [Gou03, GOST01, GOT03a]. Interior-Point [Gou03, GOT03a]. International [DMW06, DGDG97b]. interval [CG86]. introduction [CGT90a, CGT91d, GL03, SHG04]. inverse [GS98]. issue [GLT04b]. Issues [GR10b]. Iterated [CGST94b, CGST96c]. Iterated-Subspace [CGST94b, CGST96c]. iteration [CGT92h, CGT92i, CGT97a]. iterations [CGT92h, CGT92i, CGT97a]. Iterative [DGSW06, DGDG97b, Gou99a, Gou00, DGDG97a, GT02a].

J [CGT89a]. J.D [BGG15]. July [Ame94]. June [DMW06, DGDG97b].

Krylov [CGL20, GOR14].

Lagrangian [CGT97c, CGT91e, CGT92c, CGST93b, CGST93a, CGST93c, CGST96a, CGT97b, CGJR16]. LANCELOT [BCG^{+}97a, BCG^{+}97b, CGT91a, CGT96, GOT02, CGT92j, CGT90a, CGT91b, CGT92a, CGT92d, CGT92g, CGT92e]. Lanczos [GLRT99]. Large [CGT89b, CGT90b, CGT92j, CGT92f, CGT93, CGT94d, DLG94, ACD^{+}93, BCG^{+}97a, BCG^{+}97b, CGT89c, CGT90a, CGT90c, CGT92g, CGT92e, CGT94c, CGT96, DDLG96, DLG97, DDLG97, GR89, Gou91a, Gou99b, GT00, GT02a, GT02b, GS03, GOT03c, GS04, GLT04b, GHS05a, GHS05b, GOT05, GSH07, GOR13]. Large-Scale [CGT92j, CGT92f, CGT93, CGT94d, ACD^{+}93, BCG^{+}97a, BCG^{+}97b, CGT89c, CGT90c, CGT92g, CGT92e, CGT94c, CGT96, DDLG96, Gou91a, Gou99b, GT00, GT02a, GT02b, GOT03c, GOT05, GOR13]. Leading [GS10, GS09]. Least [CGT14b, CGT15, BCG^{+}10, CGT09, CGT13b, GLT04a, GT07]. Least-Squares [CGT14b, CGT15, BCG^{+}10, CGT09, CGT13b, GLT04a, GT07]. library [GOT03c]. Linear [BGNW05, CGST93a, CGST93c, DLG96, DDG98, DGDG97b, ADD^{+}94, BGG15, BGNW03, BGNW04, CGT09, CGST93b, CGST96a, CGT00a, DLG95, DDG99, DGDG97a, Gou82, Gou83a, Gou84a, Gou84b, GR89, Gou99a, Gou00, GS03, GS04, GHS05a, GHS05b, GSH07, Gou12, KGW00]. Linear-Quadratic [BGNW05]. linearly [ACD^{+}93, Gou99b]. linesearch [GLRT98, GLRT00, GOP16]. Lipschitz [FGF13]. Local [CGST93c, GR10b, GLR15]. location [CG84]. Lyngby [DMW06].

mathematical [CGT91d, GOT15]. Mathematics [Ame94]. Matrices [GS10, ADGR90, CGT91c, DGR^{+}90, DGR^{+}91, GS09]. matrix [CG86]. mature [GT04a]. memory [GLRT98]. Method [DGLR87, GR10a, GR10b, GLR14, GLR15, BBG^{+}12, DGLR90, GGM^{+}84, Gou89, GLRT99, GT02a, GOT03a, GST05, GR12]. Methods [CGT10, CGT14b, CGT15, CGST94b, CGT97d, DGDG97b, Gou03, BGG15, CGT11a, CGT11b, CGT12c, CGT13b, CGT88b, CGT94a, CGST96c, CGT00b, CGJR16, DGDG97a, GGM^{+}82a, Gou82, Gou99a, GLRT00, Gou00, GT00, GT02b, GOT05, Gou08, Gou12, GOR13, GOR14]. Michael [BGG15]. Minimization [CGT11c, CGT12d, CGST94b, CGT88b, CGT92b, CGST94a, CGT94a, CGST96b, CGST96c, Gou89, GN98, GS98]. minimizers [CGL20]. minimizing [CGT13a, CGT00a]. MINOS [BCG^{+}97a, BCG^{+}97b]. mixture [GGM^{+}82b]. models [BGG15, CGT94a]. modified [GN98, Gou99b]. monotone [GT06]. monotonic [CGL20]. motivation [CGT11a]. MR0933734 [CGT89a]. MR2533756 [GT12]. multidimensional [GLT04a]. Multifrontal [DGLR87, CGLT94, DGLR90].

necessary [GT99]. negative [GLRT00]. Nesterov [CGT13a]. net [GM81]. Newton [CGT10, CGT91c]. no [CGT89a]. non [CGOT00, GT02b, GT06]. non-convex [CGOT00, GT02b]. non-monotone [GT06]. Nonconvex [CGT10, CGT11c, CGT12d, CGT20, CGT12a, CGT00a, GT02a, GLT04b]. Nonlinear [CGT11c, CGT14b, CGT15, CGT92j, CGT92f, CGST94b, CGT97d, DG00, GT10, GLR14, ACD^{+}93, BCG^{+}10, BCG^{+}97a, BCG^{+}97b, BGNW04, CGT13b, CGT14a, CG84, CGT89c, CGT90a, CGT90c, CGT91d, CGT92a, CGT92g, CGT92h, CGT92i, CGT92e, CGT93, CGT94d, CGST96c, CGT96, CGT97a, CGOT00, FGL^{+}02, Gou88, GT00, GOST01, GOST02, GT03a, GT03b, GT03c, GOT03c, GL03, GOT03a, GT04a, GLT04a, GOT05, GT06, GT07, GLT04b, GT12]. Nonmonotone [GLR15]. norm [CGL20, GN98, GS98]. norm-minimization [GS98]. note [BGG15, CGT13a, CGT94b, CGT94a, GT99]. number [CGT92h, CGT92i, CGT97a]. Numer [CGT89a]. Numerical [CGT92g, CGT96, Gou82, GT02b, GOT05, GLR15, BCG^{+}97a, BCG^{+}97b, CGT11a, CGT92d, CGJR16, DW97, GM81, GOT02, GS03, GS04, GHS05a, GHS05b, GSH07].

occurring [GM81]. one [CGT91c]. optimality [CGT12b]. Optimization [CGT10, CGT14b, CGT15, CGT20, CGT89b, CGT90b, CGT91e, CGT92j, CGT92f, CGST93b, CGST93a, CGST93c, CGST94b, CGT97d, DLG94, DG00, DGSW10, DGDG97b, GHN01, Gou03, GLR14, ACD^{+}93, BCG^{+}97a, BCG^{+}97b, BBG^{+}12, BGNW04, CGT11a, CGT11b, CGT12a, CGT12b, CGT12c, CGT13b, CGT14a, CGT88a, CGT89a, CGT90a, CGT92a, CGT92c, CGT92g, CGT92h, CGT92i, CGT92e, CGST93d, CGT93, CGT94c, CGT94d, CGLT94, CGST96a, CGST96c, CGT96, CGT97b, CGT97a, CGT97c, DDLG96, DLG97, DDLG97, DGDG97a, FGF13, GLRT98, Gou99a, Gou99b, GLRT00, Gou00, GOT03c, GL03, GOT03a, GT04a, GLT04b, GST05, GOT05, GRT10, GOT15]. Oracle [CGT12d]. Order [CGT12d, CGT15, CGT20, BBG^{+}12, CGT12b, CGT14a, CGT94a, GT99]. other [CGT09, GRT10]. outer [CGT92h, CGT92i, CGT97a].

Package [CGT92j, CGT92a, CGT92g, CGT92e, CGT96, GT03a, GT07]. packages [BCG^{+}97a, BCG^{+}97b, GS03, GOT03c, GS04]. paper [CGT89a]. Papers [DMW06]. PARA [DMW06]. Parallel [ADD^{+}94, DMW06, DGLR87, DGLR90]. parameter [GPT12]. parameters [GOST05]. parametric [CG86]. part [DGDG97b, Gou84a, Gou84b, CGT11a, CGT11b]. Partially [DLG94, CGT94c, CGLT94, DDLG96, DLG97, DDLG97]. penalty [CG87, Gou86, Gou88, Gou89, GOT03a, GT10, GT12]. Performance [CGLT94]. Pivoting [Gou91b]. Point [DGSW06, DGSW10, Gou03, GS10, DGSW05, DGSW07, GOST01, GOT03a, GS09, GOR14]. points [CGT14a, GT99]. positive [CG86]. Potentially [CGT14b, CGT13b]. Powell [BGG15]. Practical [DAGR87, GR10b, CGJR16, Gou85]. practicalities [Gou84a, Gou84b]. practice [GLT04b]. Preconditioners [DLG94, DLG96, DLG95, DDLG96, DLG97, DDLG97, DDG99, DGSW05, DGW06, DGSW07, GS98]. Preconditioning [DGSW06, DGSW10, GOP16, KGW00]. predictor [GR12]. Preface [GLT04b, BGG15]. preliminary [SHG04]. Preprocessing [DLG96, GT04b]. primal [CGT00a, CGOT00, GOST01]. primal-dual [CGOT00, GOST01]. problem [GM81, Gou85, GT03b]. Problems [CGT10, CGT14b, DLG94, GHN01, CGT09, CGT13b, CGT88b, CGT89c, CGT90a, CGT90c, CGT91d, DDLG96, DLG97, DDLG97, DGSW05, DGSW07, GGM^{+}82a, GGM^{+}82b, Gou88, GT03a, GT07, Gou08, Gou12]. procedures [GR89]. proceedings [Ame94, DGDG97b]. Programming [BGNW05, CGT11c, GOT00, GHN01, BBG^{+}12, BGNW03, BGNW04, CG84, CG87, CGT89c, CGT90c, CGT91d, CGOT00, FGL^{+}02, GGM^{+}82a, GGM^{+}82b, GGM^{+}84, Gou82, Gou83a, Gou84a, Gou84b, Gou85, Gou88, Gou91a, GT00, GOST01, GT02a, GT02b, GT03c, GLT04b, GT04b, GT06, GT10, GT12, GOR13]. programs [Gou83b]. project [CGT90a]. Projected [GOR14]. projection [Gou08, Gou12]. projections [CGST93d]. Properties [CGST93b, CGST93c, CGT92b, CGST94a, CGST96a, CGST96b]. proposal [CGT89c, CGT90c].

Quadratic [BGNW05, GOT00, GHN01, BGG15, CGL20, GGM^{+}82a, GGM^{+}82b, GGM^{+}84, Gou82, Gou83b, Gou85, Gou91a, GT02a, GT02b, GT04b, GOR13]. quasi [CGT91c]. quasi-Newton [CGT91c].

Ranck [DAGR87]. Range [GGM^{+}82a, GGM^{+}82b]. Range-space [GGM^{+}82a, GGM^{+}82b]. Rank [CGT14b, CGT13b, CGT91c, GOST02]. Rank-Deficient [CGT14b, CGT13b]. Reflections [Gou03]. region [BGG15, CGT09, CGT88a, CGT89a, CGT92b, CGST93d, CGST94a, CGST96b, CGOT00, CGT00b, FGL^{+}02, GN98, GLRT99, GT03a, GT03c, GST05, GOST05, GT06, GT07, GRT10, GR12]. regularisation [CGT11a, CGT11b]. regularisations [CGT09]. regularised [GRT10]. Regularization [CGT14b, CGT12a, CGT12c, CGT13b, GPT12]. Regularized [CGT10, DGSW06, BCG^{+}10, CGL20, DGSW05, DGSW07]. Related [DG00]. Release [CGT92a, CGT92j, CGT92e, CGT92d, CGT92g, CGT96]. Relevance [CGT14b, CGT13b]. residual [BCG^{+}10]. Results [GOT02, GLR15, BCG^{+}97b, CGT11a, CGT92d, GS03, GHS05a]. Revised [DMW06, CGST94a]. revisited [GOT03b]. Roger [GLT04b]. Rosenbrock [CGT13a].

Saddle [DGSW06, DGSW10, GS10, DGSW05, DGSW07, GS09, GOR14, GOP16]. Saddle-Point [DGSW06, DGSW10, DGSW05, DGSW07, GOR14]. safe [GOT03c, GOT15]. Scale [CGT89b, CGT90b, CGT92j, CGT92f, DLG94, ACD^{+}93, BCG^{+}97a, BCG^{+}97b, CGT89c, CGT90a, CGT90c, CGT92g, CGT92e, CGT93, CGT94c, CGT94d, CGT96, DDLG96, DDLG97, Gou91a, Gou99b, GT00, GT02a, GT02b, GOT03c, GLT04b, GOT05, GOR13]. scaling [DGP94]. scheme [CGLT94]. Schmidt [GGM^{+}84]. Scientific [DMW06]. SDIF [CGT91d]. search [ACD^{+}93, Gou86, GOP16]. Second [CGT15, GR10a, GR10b, CGT12b, CGT94a, GT99, GR12]. second-derivative [GR12]. Second-Order [CGT15, CGT12b, CGT94a, GT99]. seismic [GM81]. Selected [DMW06]. semi [CG86, CG87]. semi-definite [CG86]. semi-infinite [CG87]. Sensitivity [GOST05]. Separable [DLG94, CGT94c, CGLT94, DDLG96, DLG97, DDLG97]. September [DGDG97b]. sequential [Gou89]. Set [Gou03, GT02a]. Sharp [CGT20]. SIAM [CGT89a]. SifDec [GOT03b]. Simple [CGT91e, CGT88a, CGT88b, CGT89a, CGT92c, CGT92h, CGT92i, CGT97b, CGT97a, CGT97c, Gou86, Gou88]. slack [CGT94b]. Smooth [CGT12d, CGT13a]. Solution [DLG95, DLG96, DDLG96, DDLG97, GHN01, GM81, Gou83b, GOST02, GS03, GS04, GHS05a, GHS05b, GSH07]. solutions [Gou85]. Solve [DLG94]. Solvers [DGSW06, GHS05a, GHS05b, GSH07, SHG04]. Solving [DDG98, GLRT99, BDG94, CGT88b, Gou88, GT03a, GT07, GRT10]. Some [Gou03]. space [GGM^{+}82a, GGM^{+}82b]. spaces [CGL20]. Sparse [DLG96, GS98, ADGR90, DGR^{+}90, DGR^{+}91, GS03, GS04, GHS05a, GHS05b, GSH07, SHG04]. Special [GLT04b]. Spectral [GS09, GS10]. SQP [FGL^{+}02, GT00, GT03c, GR10a, GR10b, GR12, GLR15]. SQP-filter [FGL^{+}02, GT03c]. Squares [CGT14b, CGT15, BCG^{+}10, CGT09, CGT13b, GLT04a, GT07]. stability [Gou83b]. standard [CGT89c, CGT90c, CGT91d]. State [DMW06, Gou03, DW97]. Steepest [CGT10, Gou83a, Gou84a, Gou84b]. Step [GLR14, GR12]. Stretching [DDG98]. structure [CGT90a, CGT94b]. structured [ADD^{+}94, CGT92b, CGST94a, CGST96b, DLG95, DDG99]. subject [CGT00a]. subproblem [GLRT99]. subproblems [BGNW04, CGT94a, GRT10]. Subspace [CGST94b, DDG99, CGST96c]. Subspace-by-subspace [DDG99]. Successive [BGNW05, BGNW03]. Superlinear [GOST01]. Supplement [CGT97c]. survey [CGT94d]. symmetric [CGT91c, DGR^{+}90, DGR^{+}91, GS03, GS04, GHS05a, GHS05b, GSH07, SHG04].

Systems [DLG96, DDG98, DGSW06, DGSW10, ADD^{+}94, BDG94, DLG95, DDG99, DGP94, GR89, Gou99a, Gou00, GOST02, GS03, GS04, GHS05a, GHS05b, GSH07, GOR14, KGW00].

techniques [GS98]. Technology [Ame94]. Testing [BCGT93, BCGT95, CGT88b, GOT03b, GOT15]. tests [CGT92d]. their [GOST05]. theory [GLT04b]. thread [GOT03c]. thread-safe [GOT03c]. threads [GOT15]. three [Gou88]. Topics [DG00]. topology [BBG^{+}12]. Toulouse [DGDG97b]. Trajectory [GOR13]. Trajectory-following [GOR13]. Trust [CGT09, CGT00b, BGG15, CGT88a, CGT89a, CGT92b, CGST93d, CGST94a, CGST96b, CGOT00, FGL^{+}02, GN98, GLRT99, GT03a, GT03c, GST05, GOST05, GT06, GT07, GRT10, GR12]. Trust-region [CGT09, CGT00b, CGOT00, FGL^{+}02, GN98, GLRT99, GT03c, GOST05, GT06, GRT10]. trust-region-free [GR12]. two [CGST93b, CGST93a, CGST93c]. type [GOP16].

Unassembled [DLG96, DDG98]. Unconstrained [BCGT93, BCGT95, CGT10, CGT11a, CGT11b, CGT12b, CGT12c, DDLG97, GLRT98, GLRT00, GOT03b, GST05, GOT15]. Unified [GLR14]. uniqueness [Gou85]. update [CGT91c]. Updating [GPT12]. USA [Ame94]. Use [ADGR90, DLG94, DDG98, DAGR87]. Using [CGT15, DLG96, DGSW05, DGSW07, BBG^{+}12, BGNW04, CGT92b, CGST93d, CGST94a, CGT94b, CGT94a, CGST96b, DLG95, DDLG96, DDLG97, GS98, GLRT99].

value [GN98]. variables

[CGT88b, CGT94b]. via [Gou88].

Weighted [GGM^{+}84]. without [GT10, GT12]. working [GT02a]. working-set [GT02a]. Workshop [DMW06, DGDG97b]. World [Ame94]. Worst [CGT20, CGT11b]. Worst-Case [CGT20, CGT11b].

Year [DGDG97b].

References

Arioli:1993:CSD

[ACD^{+}93] M. Arioli, T. F. Chan, I. S.

Duff, N. I. M. Gould, and J. K. Reid. Computing a search direction for large-scale linearly constrained nonlinear optimization calculations. Technical Report RAL-93-066 and TR-PA-93-34, Rutherford Appleton Laboratory and CERFACS, Chilton, Oxon, England and Toulouse, France, 1993. URL ftp:/

/130.246.8.32/pub/reports/
acdgrRAL93066.ps.Z.

Amestoy:1994:PAS

[ADD^{+}94] Patrick R. Amestoy, Michel J.

Daydé, Iain S. Duff, Jean-Yves L’Excellent, Nick Gould, and Chiara Puglisi. Parallel algorithms for structured systems of linear equations. In Ames [Ame94], pages 1058–1061. ISBN ???? LCCN ???? Three volumes: Volume 1: Applied mathematics, stochastic processes, iterative methods. Volume 2: Applications of computing, fluid mechanics, acoustics, solid mechanics, neural networks. Volume 3: Numerical methods and analysis finite differences and elements, differential equations, intervals.

Arioli:1990:UAC

[ADGR90] Mario Arioli, Iain S. Duff, Nicholas I. M. Gould, and John K. Reid. Use of the P^{4} and P^{5} algorithms for in-core factorization of sparse matrices. SIAM Journal on Scientific and Statistical Computing, 11(5):913–927, September 1990. CODEN SI-JCD4. ISSN 0196-5204.

Ames:1994:IPI

[Ame94] W. F. Ames, editor. IMACS ’94: proceedings of the 14th IMACS World Congress on Computation and Applied Mathematics: July 11–15, 1994, Georgia Institute of Technology, Atlanta, Georgia, USA. International Association for Mathematics and Computers in Simulation, ????, 1994. ISBN ???? LCCN ???? Three volumes: Volume 1: Applied mathematics, stochastic processes, iterative methods. Volume 2: Applications of computing, fluid mechanics, acoustics, solid mechanics, neural networks. Volume 3: Numerical methods and analysis finite differences and elements, differential equations, intervals.

Browne:2012:FMB

[BBG^{+}12] P. A. Browne, C. Budd, N. I. M. Gould, H. A. Kim, and J. A. Scott. A fast method for binary programming using firstorder derivatives, with application to topology optimization with buckling constraints. International Journal for Numerical Methods in Engineering, 92(12): 1026–1043, 2012. CODEN IJN-MBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Bongartz:1997:NCBa

[BCG^{+}97a] I. Bongartz, A. R. Conn, N. I. M.

Gould, M. A. Saunders, and Ph. L. Toint. A numerical comparison between the LANCELOT and MINOS packages for large-scale nonlinear optimization. Technical Report RAL-TR 97-054, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1997.

Bongartz:1997:NCBb

[BCG^{+}97b] I. Bongartz, A. R. Conn, N. I. M.

Gould, M. A. Saunders, and Ph. L. Toint. A numerical comparison between the LANCELOT and MINOS packages for large-scale nonlinear optimization: the complete results. Technical Report 97/14, Department of Mathematics, FUNDP, Namur, Belgium, 1997.

Bellavia:2010:CRE

[BCG^{+}10] S. Bellavia, C. Cartis, N. I. M.

Gould, B. Morini, and Ph. L. Toint. Convergence of a regularized Euclidean residual algorithm for nonlinear least-squares. SIAM Journal on Numerical Analysis, 48(1):1–29, ???? 2010. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Bongartz:1993:SCC

[BCGT93] I. Bongartz, A. R. Conn, N. I. M. Gould, and Ph. L. Toint. CUTE: Constrained and Unconstrained Testing Environment. Technical Report TR/PA/93/10, CER-FACS, Toulouse, France, 1993.

                                        URL  ftp://130.246.8.32/
                                        pub/reports/cgtCERFACS9310.

ps.Z. To appear in ACM Transactions on Mathematical Software.

Bongartz:1995:CCU

[BCGT95] I. Bongartz, A. R. Conn, Nick Gould, and Ph.L. Toint. CUTE: Constrained and unconstrained testing environment. ACM Transactions on Mathematical Software, 21(1):123–160, March 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

                                        www.acm.org/pubs/citations/
                                        journals/toms/1995-21-1/p123-
                                        bongartz/.

Bouaricha:1994:ESA

[BDG94] A. Bouaricha, Iain S. Duff, and N. I. M. Gould. Experiments in solving augmented systems. Technical Report (to appear), CERFACS, Toulouse, France, 1994.

Bienstock:2015:NFT

[BGG15] Daniel Bienstock, Philip E. Gill, and Nick Gould. A note on “On fast trust region methods for quadratic models with linear constraints”, by Michael J.D. Powell [Preface]. Mathematical Programming Computation, 7 (3):235, 2015. ISSN 1867-2949 (print), 1867-2957 (electronic).

Byrd:2003:CSL

[BGNW03] R. H. Byrd, N. I. M. Gould, J. Nocedal, and R. A. Waltz. On the convergence of successive linear programming algorithms. Technical Report RAL-TR-2003-013, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2003.

Byrd:2004:ANO

[BGNW04] Richard H. Byrd, Nicholas I. M.

Gould, Jorge Nocedal, and Richard A. Waltz. An algorithm for nonlinear optimization using linear programming and equality constrained subproblems. Mathematical Programming, 100(1, Ser. B):27– 48, 2004. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Byrd:2005:CSL

[BGNW05] Richard H. Byrd, Nicholas I. M.

Gould, Jorge Nocedal, and Richard A. Waltz. On the convergence of successive linearquadratic programming algorithms. SIAM Journal on Optimization, 16(2):471–489, ???? 2005. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/42653.

Conn:1984:LDI

[CG84] Andrew R. Conn and Nicholas I. M. Gould. On the location of directions of infinite descent for nonlinear programming algorithms. SIAM Journal on Numerical Analysis, 21(6):1162– 1179, 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Caron:1986:FPS

[CG86] R. J. Caron and N. I. M. Gould. Finding a positive semi-definite interval for a parametric matrix. Linear Algebra and its Applications, 76:19–29, 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Conn:1987:EPF

[CG87] Andrew R. Conn and Nicholas I. M. Gould. An exact penalty function for semi-infinite programming. Mathematical Programming, 37(1):19–40, 1987. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Curtis:2016:AAL

[CGJR16] Frank E. Curtis, Nicholas I. M. Gould, Hao Jiang, and Daniel P. Robinson. Adaptive augmented Lagrangian methods: algorithms and practical numerical experience. Optimization Methods & Software, 31(1):157–186, 2016. ISSN 1055-6788.

Cartis:2020:MEN

[CGL20] Coralia Cartis, Nicholas I. M. Gould, and Marius Lange. On monotonic estimates of the norm of the minimizers of regularized quadratic functions in Krylov spaces. BIT Numerical Mathematics, 60(3):??, September 2020. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL https://link.springer.
com/article/10.1007/s10543-
019-00791-2.

Conn:1994:PMS

[CGLT94] A. R. Conn, Nick I. M. Gould, M. Lescrenier, and Ph. L. Toint. Performance of a multifrontal scheme for partially separable optimization. In S. Gomez, J. P. Hennart, and R. A. Tapia, editors, Advances in optimization and numerical analysis, Proceedings of the Sixth workshop on Optimization and Numerical Analysis, Oaxaca, Mexico, 1992, volume 275 of Mathematics and its Applications Series, pages 79– 96. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994.

Conn:2000:PDT

[CGOT00] Andrew R. Conn, Nicholas I. M.

Gould, Dominique Orban, and Philippe L. Toint. A primal-dual trust-region algorithm for nonconvex nonlinear programming. Mathematical Programming, 87 (2, Ser. B):215–249, 2000. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). Studies in algorithmic optimization.

Conn:1993:GCT

[CGST93a] A. R. Conn, N. I. M. Gould, A. Sartenaer, and Ph. L. Toint.

Global convergence of two augmented Lagrangian algorithms for optimization with a combination of general equality and linear constraints. Technical Report TR/PA/93/26, CERFACS, Toulouse, France, 1993.

Conn:1993:GCP

[CGST93b] A. R. Conn, N. I. M. Gould, A. Sartenaer, and Ph. L. Toint. Global convergence properties of two augmented Lagrangian algorithms for optimization with a combination of general equality and linear constraints. Technical Report TR/PA/93/26, CER-FACS, Toulouse, France, 1993.

                                        URL  ftp://130.246.8.32/
                                        pub/reports/cgstCERFACS9326.
                                        ps.Z.

Conn:1993:LCP

[CGST93c] A. R. Conn, N. I. M. Gould, A. Sartenaer, and Ph. L. Toint. Local convergence properties of two augmented Lagrangian algorithms for optimization with a combination of general equality and linear constraints. Technical Report TR/PA/93/27, CER-FACS, Toulouse, France, 1993.

                                        URL  ftp://130.246.8.32/
                                        pub/reports/cgstCERFACS9327.
                                        ps.Z.

Conn:1993:GCC

[CGST93d] A. R. Conn, Nick I. M. Gould, A. Sartenaer, and Ph. L. Toint. Global convergence of a class of trust region algorithms for optimization using inexact projections on convex constraints.

SIAM Journal on Optimization, 3(1):164–221, 1993. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Conn:1994:CPM

[CGST94a] A. R. Conn, N. I. M. Gould, A. Sartenaer, and Ph. L. Toint. Convergence properties of minimization algorithm for convex constraints using a structured trust region (revised). Technical Report 94/5, Department of Mathematics, FUNDP, Namur, Belgium, 1994. URL ftp:

//138.48.4.14/pub/reports/
TR94-05.ps.

Conn:1994:ISM

[CGST94b] A. R. Conn, N. I. M. Gould, A. Sartenaer, and Ph. L. Toint. On iterated-subspace minimization methods for nonlinear optimization. Technical Report 94-069, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1994. URL ftp:/

/130.246.8.32/pub/reports/
cgstRAL94069.ps.Z.

Conn:1996:CPA

[CGST96a] A. R. Conn, N. Gould, A. Sartenaer, and Ph. L. Toint. Convergence properties of an augmented Lagrangian algorithm for optimization with a combination of general equality and linear constraints. SIAM Journal on Optimization, 6(3):674– 703, 1996. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/25146.

Conn:1996:CPM

[CGST96b] A. R. Conn, Nick Gould, A. Sartenaer, and Ph. L. Toint. Convergence properties of minimization algorithms for convex constraints using a structured trust region. SIAM Journal on Optimization, 6(4):1059– 1086, 1996. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/23648.

Conn:1996:ISM

[CGST96c] A. R. Conn, Nick Gould, A. Sartenaer, and Ph. L. Toint. On iterated-subspace minimization methods for nonlinear optimization. In Linear and nonlinear conjugate gradientrelated methods (Seattle, WA, 1995), pages 50–78. SIAM Press, Philadelphia, PA, USA, 1996.

Conn:1988:GCC

[CGT88a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Global convergence of a class of trust region algorithms for optimization with simple bounds. SIAM Journal on Numerical Analysis, 25(2):433– 460, 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). See also same journal 26(3), 764–767, 1989.

Conn:1988:TCM

[CGT88b] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint.

Testing a class of methods for solving minimization problems with simple bounds on the variables. Mathematics of Computation, 50(182):399–430, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Conn:1989:CPG

[CGT89a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Correction to the paper: “Global convergence of a class of trust region algorithms for optimization with simple bounds” [SIAM J. Numer. Anal. 25 (1988), no. 2, 433–460; MR0933734 (89h:90192)]. SIAM Journal on Numerical Analysis, 26(3):764– 767, 1989. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1989:LSO

[CGT89b] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Large scale optimization. Mathematical Programming, Series B, 45(3), 1989. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Conn:1989:PSD

[CGT89c] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. A proposal for a standard data input format for large-scale nonlinear programming problems. Report CSS-89-61, University of Waterloo, Waterloo, ON, Canada, 1989.

Conn:1990:ISL

[CGT90a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. An introduction to the structure of large scale nonlinear optimization problems and the LANCELOT project. In R. Glowinski and A. Lichnewsky, editors, Computing Methods in Applied Sciences and Engineering (Paris, 1990), pages 42–51. SIAM Press, Philadelphia, PA, USA, 1990.

Conn:1990:LSO

[CGT90b] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Large scale optimization — applications. Mathematical Programming, Series B, 48(1–3), March 1990. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Conn:1990:PSD

[CGT90c] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. A proposal for a standard data input format for large-scale nonlinear programming problems. Report CS-89-61, Computer Science Department, University of Waterloo, Ontario, Canada, 1990.

Conn:1990:F

[CGT90d] A. R. Conn, Nick Gould, and P. L. Toint. Foreword. Mathematical Programming, 48(1, (Ser. B)):vii, 1990. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Conn:1991:CDL

[CGT91a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. A comprehensive description of LANCELOT. Technical Report 91/10, Department of Mathematics, FUNDP, Namur, Belgium, 1991.

Conn:1991:CDS

[CGT91b] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. A comprehensive description of LANCELOT. Technical Report 91/10, Department of Mathematics, FUNDP, Namur, Belgium, 1991.

Conn:1991:CQN

[CGT91c] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Convergence of quasi-Newton matrices generated by the symmetric rank one update. Mathematical Programming, 50(2, (Ser. A)):177– 195, 1991. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Conn:1991:ISD

[CGT91d] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. An introduction the Standard Data Input Format (SDIF) for nonlinear mathematical programming problems. Technical Report 91/8, Department of Mathematics, FUNDP, Namur, Belgium, 1991.

Conn:1991:GCA

[CGT91e] Andrew R. Conn, Nicholas I. M.

Gould, and Philippe L. Toint. A globally convergent augmented Lagrangian algorithm for optimization with general constraints and simple bounds. SIAM Journal on Numerical Analysis, 28(2):545–572, 1991.

CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Conn:1992:ARS

[CGT92a] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Announcement of the release A of LANCELOT, a package for nonlinear optimization. Technical Report 92/4, Department of Mathematics, FUNDP, Namur, Belgium, 1992.

                                        URL http://www.rl.ac.uk/
                                        departments/ccd/numerical/
                                        lancelot/blurb.html.

Conn:1992:CPM

[CGT92b] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Convergence properties of minimization algorithms for convex constraints using a structured trust region. Technical Report 92-069, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1992. URL ftp:

                                        //130.246.8.32/pub/reports/
                                        cgtRAL92069.ps.Z.

Conn:1992:GCL

[CGT92c] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. A globally convergent Lagrangian barrier algorithm for optimization with general inequality constraints and simple bounds. Technical Report 92-067, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1992. URL ftp:/

                                        /130.246.8.32/pub/reports/
                                        cgtRAL92067.ps.Z.

Conn:1992:INT

[CGT92d] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Intensive numerical tests with LANCELOT (Release A): the complete results. Technical Report 92/15, Department of Mathematics, FUNDP, Namur, Belgium, 1992. URL ftp:

//130.246.8.32/pub/reports/
cgtFUNDP9215.ps.Z.

Conn:1992:SLF

[CGT92e] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. LANCELOT: a Fortran package for large-scale nonlinear optimization (Release A). Number 17 in Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55470-X.

Conn:1992:LSN

[CGT92f] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Large-scale nonlinear constrained optimization. In R. E. O’Malley, Jr., editor, Proceedings of the Second International Conference on Industrial and Applied Mathematics (ICIAM 91, Washington, DC, 1991), pages 51–70. SIAM Press, Philadelphia, PA, USA, 1992.

Conn:1992:NES

[CGT92g] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. Numerical experiments with the LANCELOT package (Release A) for large-scale nonlinear optimization. Technical Report 92-075, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1992. URL ftp:/

                                        /130.246.8.32/pub/reports/
                                        cgtFUNDP9216.ps.Z.

Conn:1992:NIIa

[CGT92h] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. On the number of inner iterations per outer iteration of a globally convergent algorithm for optimization with general nonlinear equality constraints and simple bounds. In D. F. Griffiths and G. A. Watson, editors, Numerical analysis 1991 (Dundee, 1991), volume 260 of Pitman Research Notes in Mathematics Series, pages 49– 68. Longmans, Harlow, Essex, England, 1992.

Conn:1992:NIIb

[CGT92i] A. R. Conn, N. I. M. Gould, and Ph. L. Toint. On the number of inner iterations per outer iteration of a globally convergent algorithm for optimization with general nonlinear inequality constraints and simple bounds. Technical Report 92-068, Rutherford Appleton Laboratory, Chilton, Oxon, England,

                                        1992. URL ftp://130.246.8.
                                        32/pub/reports/cgtRAL92068.
                                        ps.Z.

Conn:1992:LFP

[CGT92j] Andrew R. Conn, Nick I. M. Gould, and Philippe L. Toint. Lancelot: A Fortran Package for Large-Scale Nonlinear Optimization (Release A). Number 17 in Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1992. ISBN 3-540-55470-X (Berlin, Heidelberg), 0-387-55470-X (New York). xviii + 330 pp. LCCN QA402.5 .C648 1992. US$89.00; US$118.95. URL http:/

/www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
038755470X.

Conn:1993:LSN

[CGT93] A. R. Conn, N. Gould, and Ph. L. Toint. Large-scale nonlinear constrained optimization. In Linear algebra for large scale and real-time applications (Leuven, 1992), volume 232 of NATO Adv. Sci. Inst. Ser. E Appl. Sci., pages 21–48. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993.

Conn:1994:NUA

[CGT94a] A. R. Conn, Nick I. M. Gould, and Ph. L. Toint. A note on using alternative secondorder models for the subproblems arising in barrier function methods for minimization. Numerische Mathematik, 68(1): 17–33, June 1994. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://link.springer.
de/link/service/journals/
00211/bibs/4068001/40680017.
htm;   http://science.springer.
de/nmee/bibs/4068001/40680017.
htm.

Conn:1994:NES

[CGT94b] Andrew R. Conn, Nick I. M.

Gould, and Ph. L. Toint. A note on exploiting structure when using slack variables. Mathematical Programming, 67(1–3):89– 97, 1994. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL http:/

                                        /link.springer.com/article/
                                        10.1007/BF01582214.

Conn:1994:IDP

[CGT94c] Andrew R. Conn, Nick I. M. Gould, and Philippe L. Toint. Improving the decomposition of partially separable functions in the context of large-scale optimization: a first approach. In W. W. Hager, D. W. Hearn, and P. M. Pardalos, editors, Large scale optimization: State of the Art (Gainesville, FL, 1993), pages 82–94. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994.

Conn:1994:LSN

[CGT94d] Andrew R. Conn, Nick I. M. Gould, and Philippe L. Toint. Large-scale nonlinear constrained optimization: a current survey. In E. Spedicato, editor, Algorithms for continuous optimization: the state of the art (Il Ciocco, 1993), volume 434 of NATO Adv. Sci. Inst. Series C: Mathematical and Physical Sciences, pages 287–332. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1994.

Conn:1996:NEL

[CGT96] A. R. Conn, Nick Gould, and Ph. L. Toint. Numerical experiments with the LANCELOT package (Release A) for largescale nonlinear optimization. Mathematical Programming, 73 (1, Ser. A):73–110, 1996. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Conn:1997:NII

[CGT97a] A. R. Conn, N. Gould, and Ph. L. Toint. On the number of inner iterations per outer iteration of a globally convergent algorithm for optimization with general nonlinear inequality constraints and simple bounds. Computational optimization and applications, 7(1):41–69, 1997. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic). Computational issues in high performance software for nonlinear optimization (Capri, 1995).

Conn:1997:GCL

[CGT97b] A. R. Conn, Nick Gould, and Ph. L. Toint. A globally convergent Lagrangian barrier algorithm for optimization with general inequality constraints and simple bounds. Mathematics of Computation, 66 (217):261–288, S1–S11, January 1997. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/
jour-pbprocess?fn=110&arg1=
S0025-5718-97-00777-1&u=/
mcom/1997-66-217/.

Conn:1997:SGC

[CGT97c] A. R. Conn, Nick Gould, and Ph. L. Toint. Supplement to “A globally convergent Lagrangian barrier algorithm for optimization with general inequality constraints and simple bounds”. Mathematics of Computation, 66(217):S1–S11, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Conn:1997:MNC

[CGT97d] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. Methods for nonlinear constraints in optimization calculations. In Duff and Watson [DW97], pages 363–390. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

                                        //global.oup.com/academic/
                                        product/the-state-of-the-
                                        art-in-numerical-analysis-
                                        9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Conn:2000:PDA

[CGT00a] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. A primal–dual algorithm for minimizing a nonconvex function subject to bound and linear equality constraints. In Di Pillo and Giannessi [DG00], pages 15–50. ISBN 0-7923-6109-1 (hardcover).

LCCN QA402.5 .N55 2000.

URL  http://www.loc.gov/
catdir/enhancements/fy0822/
99056953-d.html;   http://
www.loc.gov/catdir/enhancements/
fy0822/99056953-t.html.

Conn:2000:TRM

[CGT00b] Andrew R. Conn, Nicholas I. M. Gould, and Philippe L. Toint. Trust-region methods. MPS/SIAM Series on Optimization. SIAM Press, Philadelphia, PA, USA, 2000. ISBN 0-89871-460-5. xx + 959 pp.

Cartis:2009:TRO

[CGT09] C. Cartis, N. I. M. Gould, and P. L. Toint. Trust-region and other regularisations of linear least-squares problems. BIT Numerical Mathematics, 49(1):21– 53, March 2009. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=49&issue=1&spage=
21.

Cartis:2010:CSD

[CGT10] C. Cartis, N. I. M. Gould, and Ph. L. Toint. On the complexity of steepest descent, Newton’s and regularized Newton’s methods for nonconvex unconstrained optimization problems. SIAM Journal on Optimization, 20(6): 2833–2852, ???? 2010. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Cartis:2011:ACRa

[CGT11a] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. Adaptive cubic regularisation methods for unconstrained optimization. Part I: motivation, convergence and numerical results. Mathematical Programming, 127(2, Ser. A):245–295, 2011. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Cartis:2011:ACRb

[CGT11b] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. Adaptive cubic regularisation methods for unconstrained optimization. Part II: worstcase function- and derivativeevaluation complexity. Mathematical Programming, 130(2, Ser. A):295–319, 2011. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Cartis:2011:ECC

[CGT11c] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. On the evaluation complexity of composite function minimization with applications to nonconvex nonlinear programming. SIAM Journal on Optimization, 21(4):1721–1739, ???? 2011. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/siopt/resource/1/
                                        sjope8/v21/i4/p1721_s1.

Cartis:2012:ACR

[CGT12a] C. Cartis, N. I. M. Gould, and Ph. L. Toint. An adaptive cubic regularization algorithm for nonconvex optimization with convex constraints and its functionevaluation complexity. IMA Journal of Numerical Analysis, 32(4):1662–1695, October 2012. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (elec-

tronic). URL http://imajna.
oxfordjournals.org/content/
32/4/1662.full.pdf+html.

Cartis:2012:CBS

[CGT12b] C. Cartis, N. I. M. Gould, and Ph. L. Toint. Complexity bounds for second-order optimality in unconstrained optimization. Journal of Complexity, 28 (1):93–108, 2012. CODEN JO-COEH. ISSN 0885-064x (print), 1090-2708 (electronic).

Cartis:2012:ECA

[CGT12c] Coralia Cartis, Nicholas I. M.

Gould, and Philippe L. Toint. Evaluation complexity of adaptive cubic regularization methods for convex unconstrained optimization. Optimization Methods & Software, 27(2):197–219, 2012. ISSN 1055-6788.

Cartis:2012:OCF

[CGT12d] Coralia Cartis, Nicholas I. M.

Gould, and Philippe L. Toint. On the oracle complexity of first-order and derivative-free algorithms for smooth nonconvex minimization. SIAM Journal on Optimization, 22 (1):66–86, ???? 2012. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/siopt/resource/1/
                                        sjope8/v22/i1/p66_s1.

Cartis:2013:NAC

[CGT13a] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. A note about the complexity of minimizing Nesterov’s smooth Chebyshev–Rosenbrock function. Optimization Methods & Software, 28(3):451–457, 2013. ISSN 1055-6788.

Cartis:2013:ECC

[CGT13b] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. On the evaluation complexity of cubic regularization methods for potentially rank-deficient nonlinear least-squares problems and its relevance to constrained nonlinear optimization. SIAM Journal on Optimization, 23(3):1553– 1574, 2013. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Cartis:2014:CFF

[CGT14a] Coralia Cartis, Nicholas I. M. Gould, and Philippe L. Toint. On the complexity of finding first-order critical points in constrained nonlinear optimization. Mathematical Programming, 144 (1-2, Ser. A):93–106, 2014. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Cartis:2014:ECC

[CGT14b] Coralia Cartis, Nicholas I. M.

Gould, and Philippe L. Toint. On the evaluation complexity of cubic regularization methods for potentially rank-deficient nonlinear least-squares problems and its relevance to constrained nonlinear optimization. SIAM Journal on Optimization, 23(3): 1553–1574, ???? 2014. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Cartis:2015:ECC

[CGT15] Coralia Cartis, Nicholas I. M.

Gould, and Philippe L. Toint. On the evaluation complexity of constrained nonlinear least-squares and general constrained nonlinear optimization using secondorder methods. SIAM Journal on Numerical Analysis, 53(2): 836–851, ???? 2015. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cartis:2020:SWC

[CGT20] Coralia Cartis, Nicholas I. M.

Gould, and Philippe L. Toint. Sharp worst-case evaluation complexity bounds for arbitraryorder nonconvex optimization with inexpensive constraints. SIAM Journal on Optimization, 30(1):513–541, ???? 2020. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Duff:1987:PUH

[DAGR87] I. S. Duff, M. Anoli, N. I. M.

Gould, and J. K. Reid. The practical use of the Hellerman–Ranck P^{4} algorithm and the P^{5} algorithm of Erisman et al. Report CSS213, AERE Harwell Laboratory, Chilton, Oxon, England, 1987.

Dayde:1998:UBS

[DDG98] M. J. Daydé, J. P. Décamps, and N. I. M. Gould. On the use of block stretching for solving unassembled linear systems. Calculateurs Paralléles, Réseaux et Systémes Répartis, 10(4):391– 399, ???? 1998. ISSN 1260-3198.

Dayde:1999:SSP

[DDG99] Michel J. Daydé, Jérôme P. Décamps, and Nicholas I. M. Gould. Subspace-by-subspace preconditioners for structured linear systems. Numerical Linear Algebra with Applications, 6(3): 213–234, April/May 1999. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Dayde:1996:SLS

[DDLG96] Michel J. Daydé, Jérôme P. Décamps, Jean-Yves L’Excellent, and Nicholas I. M. Gould. Solution of large-scale partially separable optimization problems using element-by-element preconditioners. In Poster Session of the Fifth SIAM Conference on Optimization. SIAM Press, Philadelphia, PA, USA, 1996.

Dayde:1997:SLS

[DDLG97] M. J. Daydé, J. P. Décamps, J.-Y. L’Excellent, and N. I. M. Gould. Solution of large scale partially separable unconstrained optimization problems using element-by-element preconditioners. In ????, editor, Proceedings of NAFEMS World Congress 97, Glasgow, Scotland, volume 2, pages 942–953. ????, ????, 1997. URL http://www.

nafems.org/.

DiPillo:2000:NOR

[DG00] Gianni Di Pillo and Franco Giannessi, editors. Nonlinear Optimization and Related Topics, Applied optimization. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6109-1 (hardcover). LCCN QA402.5 .N55 2000.

URL  http://www.loc.gov/
catdir/enhancements/fy0822/
99056953-d.html;   http://
www.loc.gov/catdir/enhancements/
fy0822/99056953-t.html.

Duff:1997:DML

[DGDG97a] I. S. Duff, N. I. M. Gould, C. C.

Douglas, and L. Giraud. Direct methods, linear algebra in optimization, iterative methods. BIT Numerical Mathematics, 37(3): ??, September 1997. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). Proceedings from the International Linear Algebra Year Workshops. September 1995– June 1996.

Duff:1997:PWD

[DGDG97b] Iain Duff, Nick Gould, Craig Douglas, and Luc Giraud, editors. Proceedings of the Workshop on Direct Methods, September 26–29, 1995; the Workshop on Linear Algebra in Optimization, April 22–25, 1996; and the Workshop on Iterative Methods, June 10–13, 1996, held in Toulouse as part of the International Linear Algebra Year (ILAY). BIT, Copenhagen, 1997. CODEN NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). BIT 37 (1997), no. 3.

Duff:1987:MMP

[DGLR87] I. Duff, N. Gould, M. Lescrenier, and J. Reid. The multifrontal method in a parallel environment. Technical Report CSS-211, AERE Harwell Laboratory, Chilton, Oxon, England, 1987.

Duff:1990:MMP

[DGLR90] I. S. Duff, N. I. M. Gould, M. Lescrenier, and J. K. Reid. The multifrontal method in a parallel environment. In M. G. Cox and S. J. Hammarling, editors, Reliable Scientific Computation, Oxford Sci. Publ., pages 93–111. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990.

Duff:1994:ESA

[DGP94] Iain S. Duff, N. I. M. Gould, and J. Patricio. Experiments in scaling augmented systems. Technical Report (to appear), Rutherford Appleton Laboratory, Chilton, Oxon, England, 1994.

Duff:1990:FSS

[DGR^{+}90] I. S. Duff, N. I. M. Gould, J. K. Reid, J. A. Scott, and K. Turner.

The factorization of sparse symmetric indefinite matrices. Report RAL-90-066, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1990. ???? pp.

Duff:1991:FSS

[DGR^{+}91] I. S. Duff, N. I. M. Gould, J. K.

Reid, J. A. Scott, and K. Turner. The factorization of sparse symmetric indefinite matrices. IMA Journal of Numerical Analysis, 11(2):181–204, 1991. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Dollar:2005:UCP

[DGSW05] H. S. Dollar, N. I. M. Gould, W. H. A. Schilders, and A. J. Wathen. Using constraint preconditioners with regularized saddle-point problems. Technical Report RAL-TR-2005-026, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2005.

Dollar:2006:IFP

[DGSW06] H. Sue Dollar, Nicholas I. M.

Gould, Wil H. A. Schilders, and Andrew J. Wathen. Implicitfactorization preconditioning and iterative solvers for regularized saddle-point systems. SIAM Journal on Matrix Analysis and Applications, 28(1):170– 189, January 2006. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Dollar:2007:UCP

[DGSW07] H. S. Dollar, N. I. M. Gould, W. H. A. Schilders, and A. J.

Wathen. Using constraint preconditioners with regularized saddle-point problems. Computational optimization and applications, 36(2-3):249–270, 2007. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Dollar:2010:PSP

[DGSW10] H. Sue Dollar, Nicholas I. M. Gould, Martin Stoll, and Andrew J. Wathen. Preconditioning saddle-point systems with applications in optimization. SIAM Journal on Scientific Computing, 32(1):249–270, ???? 2010. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Dollar:2006:IFC

[DGW06] H. Sue Dollar, Nicholas I. M. Gould, and Andrew J. Wathen. On implicit-factorization constraint preconditioners. In Large-scale nonlinear optimization, volume 83 of Nonconvex Optim. Appl., pages 61– 82. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Dayde:1994:UEE

[DLG94] M. Daydé, J.-Y. L’Excellent, and N. I. M. Gould. On the use of element-by-element preconditioners to solve large scale partially separable optimization problems. Technical Report RT/APO/94/4, ENSEEIHT-IRIT, Toulouse, France, 1994.

URL  ftp://130.246.8.32/

pub/reports/degENSEEIHT944.
ps.Z.

Dayd:1995:SSS

[DLG95] M. J. Daydé, J.-Y. L’Excellent, and N. I. M. Gould. Solution of structured systems of linear equations using elementby-element preconditioners. In Svetozar D. Margenov and Panayot S. Vassilevski, editors, Iterative methods in linear algebra, II. Proceedings of the 2nd IMACS international symposium on iterative methods in linear algebra, Blagoevgrad, Bulgaria, June 17–20, 1995 (1996), volume 3 of IMACS series in computational and applied mathematics, page ?? IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, 1995. ISSN 1098-870X. LCCN QA1 I617 v. 3.

Dayde:1996:PSU

[DLG96] M. J. Daydé, J.-Y. L’Excellent, and N. I. M. Gould. Preprocessing of sparse unassembled linear systems for efficient solution using element-by-element preconditioners. Lecture Notes in Computer Science, 1124:34–43, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dayde:1997:EEP

[DLG97] Michel J. Daydé, Jean-Yves L’Excellent, and Nicholas I. M. Gould. Element-by-element preconditioners for large partially separable optimization problems. SIAM Journal on Scientific Computing, 18(6):1767– 1787, November 1997. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/27479.

Dongarra:2006:APC

[DMW06] Jack Dongarra, Kaj Madsen, and Jerzy Wasniewski, editors. Applied Parallel Computing: State of the Art in Scientific Computing. 7th International Workshop, PARA 2004, Lyngby, Denmark, June 20–23, 2004. Revised Selected Papers, volume 3732 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-29067-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-29067-4;   http:/
/www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3732.

Duff:1997:SAN

[DW97] Iain S. Duff and G. Alistair Watson, editors. The state of the art in numerical analysis, volume 63 of The Institute of Mathematics and Its Applications conference series: new series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-850014-9. LCCN QA297.S775 1997. URL https:

//global.oup.com/academic/

product/the-state-of-the-
art-in-numerical-analysis-
9780198500148. Based on the

proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Fowkes:2013:BBA

[FGF13] Jaroslav M. Fowkes, Nicholas I. M. Gould, and Chris L. Farmer. A branch and bound algorithm for the global optimization of Hessian Lipschitz continuous functions. Journal of Global Optimization, 56(4):1791– 1815, 2013. CODEN JGOPEO. ISSN 0925-5001 (print), 1573-2916 (electronic).

Fletcher:2002:GCT

[FGL^{+}02] Roger Fletcher, Nicholas I. M.

Gould, Sven Leyffer, Philippe L. Toint, and Andreas Wächter. Global convergence of a trustregion SQP-filter algorithm for general nonlinear programming. SIAM Journal on Optimization, 13(3):635–659, 2002. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/35725.

Gill:1982:RSM

[GGM^{+}82a] P. E. Gill, N. I. M. Gould, W. Murray, M. A. Saunders, and M. H. Wright. Range-space methods for convex quadratic programming problems. Technical report SOL 82-9, Department of Operations Research, Stanford University, Stanford, CA, USA, 1982.

Gill:1982:RSQ

[GGM^{+}82b] P. E. Gill, N. I. M. Gould, W. Murray, M. A. Saunders, and M. H. Wright. A rangespace quadratic programming algorithm for problems with a mixture of bounds and general constraints. Technical Report SOL 82-10, Department of Operations Research, Stanford University, Stanford, CA, USA, 1982.

Gill:1984:WGS

[GGM^{+}84] Philip E. Gill, Nicholas I. M. Gould, Walter Murray, Michael A. Saunders, and Margaret H. Wright. A weighted Gram– Schmidt method for convex quadratic programming. Mathematical Programming, 30(2): 176–196, 1984. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:2001:SEC

[GHN01] Nicholas I. M. Gould, Mary E. Hribar, and Jorge Nocedal. On the solution of equality constrained quadratic programming problems arising in optimization. SIAM Journal on Scientific Computing, 23(4):1376–1395, July 2001. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

                                        epubs.siam.org/sam-bin/dbq/
                                        article/34566.

Gould:2005:CRN

[GHS05a] N. I. M. Gould, Y. Hu, and J. A.

Scott. Complete results from a numerical evaluation of sparse direct solvers for the solution of large, sparse, symmetric linear systems of equations. Numerical Analysis Group Internal Report 2005-1, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2005.

Gould:2005:NES

[GHS05b] N. I. M. Gould, Y. Hu, and J. A.

Scott. A numerical evaluation of sparse direct solvers for the solution of large, sparse, symmetric linear systems of equations. Technical Report RAL-TR-2005-005, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2005.

Gould:2003:IAN

[GL03] Nicholas I. M. Gould and Sven Leyffer. An introduction to algorithms for nonlinear optimization. In Frontiers in numerical analysis (Durham, 2002), Universitext, pages 109–197. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003.

Gould:2014:FMU

[GLR14] Nicholas I. M. Gould, Yueling Loh, and Daniel P. Robinson. A filter method with unified step computation for nonlinear optimization. SIAM Journal on Optimization, 24(1):175–209, ???? 2014. CODEN SJOPE8.

ISSN 1052-6234 (print), 1095-7189 (electronic).

Gould:2015:NFS

[GLR15] Nicholas I. M. Gould, Yueling Loh, and Daniel P. Robinson. A nonmonotone filter SQP method: Local convergence and numerical results. SIAM Journal on Optimization, 25(3):1885–1911, ???? 2015. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Gould:1998:LAM

[GLRT98] Nicholas I. M. Gould, Stefano Lucidi, Massimo Roma, and Philippe L. Toint. A linesearch algorithm with memory for unconstrained optimization. In High performance algorithms and software in nonlinear optimization (Ischia, 1997), volume 24 of Appl. Optim., pages 207–223. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

Gould:1999:STR

[GLRT99] Nicholas I. M. Gould, Stefano Lucidi, Massimo Roma, and Philippe L. Toint. Solving the trust-region subproblem using the Lanczos method. SIAM Journal on Optimization, 9(2): 504–525, March 1999. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/32273.

Gould:2000:ENC

[GLRT00] N. I. M. Gould, S. Lucidi, M. Roma, and Ph. L. Toint. Exploiting negative curvature directions in linesearch methods for unconstrained optimization. Optimization Methods & Software, 14(1-2):75–98, 2000. CODEN OMSOE2. ISSN 1055-6788. International Conference on Nonlinear Programming and Variational Inequalities (Hong Kong, 1998).

Gould:2004:MFA

[GLT04a] Nicholas I. M. Gould, Sven Leyffer, and Philippe L. Toint. A multidimensional filter algorithm for nonlinear equations and nonlinear least-squares. SIAM Journal on Optimization, 15(1):17– 38, 2004. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/42263.

Gould:2004:PNP

[GLT04b] Nick Gould, Sven Leyffer, and Philippe Toint. Preface: Nonlinear programming: theory and practice [Special issue on large scale nonconvex optimization algorithms and applications in honor of Roger Fletcher’s 65th birthday]. Mathematical Programming, 100(1, Ser. B):1– 2, 2004. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:1981:NSP

[GM81] N. I. M. Gould and W. Murray.

The numerical solution of a problem arising from the accurate determination of an earthquake epicenter occurring in a seismic net. Report, Stanford University and U.S. Geological Survey, Stanford, CA, USA, 1981.

Gould:1998:MAV

[GN98] Nicholas I. M. Gould and Jorge Nocedal. The modified absolute-value factorization norm for trust-region minimization. In High performance algorithms and software in nonlinear optimization (Ischia, 1997), volume 24 of Appl. Optim., pages 225–241. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

Gould:2016:DTS

[GOP16] N. Gould, C. Ortner, and D. Packwood. A dimertype saddle search algorithm with preconditioning and linesearch. Mathematics of Computation, 85(302):2939–2966, November 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL  http://www.ams.org/
journals/mcom/2016-85-302/
S0025-5718-2016-03096-3;
http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-
2016-03096-3/S0025-5718-2016-
03096-3.pdf;   http://www.
ams.org/mathscinet/search/
author.html?authorName=Packwood%
2C%20D.;   http://www.ams.
org/mathscinet/search/author.

html?mrauthid=75720;   http:
//www.ams.org/mathscinet/
search/author.html?mrauthid=
803698.

Gould:2013:TFM

[GOR13] Nicholas I. M. Gould, Dominique Orban, and Daniel P. Robinson. Trajectory-following methods for large-scale degenerate convex quadratic programming. Mathematical Programming Computation, 5(2):113–142, 2013. ISSN 1867-2949 (print), 1867-2957 (electronic).

Gould:2014:PKM

[GOR14] Nick Gould, Dominique Orban, and Tyrone Rees. Projected Krylov methods for saddle-point systems. SIAM Journal on Matrix Analysis and Applications, 35(4):1329–1343, 2014. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Gould:2001:SCP

[GOST01] Nicholas I. M. Gould, Dominique Orban, Annick Sartenaer, and Philippe L. Toint. Superlinear convergence of primal-dual interior point algorithms for nonlinear programming. SIAM Journal on Optimization, 11(4):974– 1002, March/May 2001. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/37051.

Gould:2002:CFC

[GOST02] Nicholas I. M. Gould, Dominique Orban, Annick Sartenaer, and Philippe L. Toint. Componentwise fast convergence in the solution of full-rank systems of nonlinear equations. Mathematical Programming, 92(3, Ser. B):481– 508, 2002. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). ISMP 2000, Part 2 (Atlanta, GA).

Gould:2005:STR

[GOST05] Nicholas I. M. Gould, Dominique Orban, Annick Sartenaer, and Phillipe L. Toint. Sensitivity of trust-region algorithms to their parameters. 4OR, 3(3):227–241, September 2005. ISSN 1619-4500.

Gould:2000:QPB

[GOT00] N. I. M. Gould, D. Orban, and Ph. L. Toint. A quadratic programming bibliography. Numerical Analysis Group Internal Report 2000-1, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2000.

Gould:2002:RNE

[GOT02] N. I. M. Gould, D. Orban, and Ph. L. Toint. Results from a numerical evaluation of LANCELOT B. Numerical Analysis Group Internal Report 2002-1, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2002.

Gould:2003:IPP

[GOT03a] N. I. M. Gould, D. Orban, and Ph. L. Toint. An interior-point l[1]-penalty method for nonlinear optimization. Technical Report RAL-TR-2003-022, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2003.

Gould:2003:CSC

[GOT03b] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. CUTEr and SifDec: A constrained and unconstrained testing environment, revisited. ACM Transactions on Mathematical Software, 29(4):373–394, December 2003. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2003:GLT

[GOT03c] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. GALAHAD, a library of threadsafe Fortran 90 packages for large-scale nonlinear optimization. ACM Transactions on Mathematical Software, 29(4): 353–372, December 2003. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2005:NML

[GOT05] Nick Gould, Dominique Orban, and Philippe Toint. Numerical methods for large-scale nonlinear optimization. Acta Numerica, 14:299–361, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Gould:2015:CUT

[GOT15] Nicholas I. M. Gould, Dominique Orban, and Philippe L. Toint. CUTEst: a constrained and unconstrained testing environment with safe threads for mathematical optimization. Computational optimization and applications, 60(3):545–557, 2015. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Gould:1982:NML

[Gou82] N. I. M. Gould. Numerical methods for linear and quadratic programming. PhD thesis, University of Oxford, Oxford, England, 1982.

Gould:1983:GSE

[Gou83a] N. I. M. Gould. The generalized steepest edge for linear programming. Research report CORR 83-2, Department of Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada, 1983.

Gould:1983:SSG

[Gou83b] N. I. M. Gould. The stability of the solution of general quadratic programs. Research report CORR 83-11, Department of Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada, 1983.

Gould:1984:GSE

[Gou84a] N. I. M. Gould. The generalized steepest–edge for linear programming: part 2, practicalities. Report CORR 84-1, Department of Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada,

Gould:1984:SEL

[Gou84b] N. I. M. Gould. The steepest edge for linear programming, part 2: practicalities. Research report CORR 84-1, Department of Combinatorics and Optimization, University of Waterloo, Ontario, Canada, 1984.

Gould:1985:PCE

[Gou85] Nicholas I. M. Gould. On practical conditions for the existence and uniqueness of solutions to the general equality quadratic programming problem. Mathematical Programming, 32(1):90– 99, 1985. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:1986:ADS

[Gou86] Nicholas Ian Mark Gould. On the accurate determination of search directions for simple differentiable penalty functions. IMA Journal of Numerical Analysis, 6(3):357–372, 1986. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Gould:1988:STC

[Gou88] N. I. M. Gould. On solving three classes of nonlinear programming problems via simple differentiable penalty functions. Journal of Optimization Theory and Applications, 56(1):89– 126, 1988. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Gould:1989:CSP

[Gou89] Nicholas Ian Mark Gould. On the convergence of a sequential penalty function method for constrained minimization. SIAM Journal on Numerical Analysis, 26(1):107–128, 1989. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Gould:1991:ALS

[Gou91a] Nicholas I. M. Gould. An algorithm for large-scale quadratic programming. IMA Journal of Numerical Analysis, 11(3):299– 324, 1991. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Gould:1991:GGE

[Gou91b] Nick I. M. Gould. On growth in Gaussian elimination with complete pivoting. SIAM Journal on Matrix Analysis and Applications, 12(2):354–361, April 1991. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Gould:1999:IMI

[Gou99a] N. I. M. Gould. Iterative methods for ill–conditioned linear systems from optimization. In Di Pillo and Giannessi [DG00], pages 123–142. ISBN 0-7923-6109-1 (hardcover). LCCN QA402.5 .N55 2000.

URL  http://www.loc.gov/
catdir/enhancements/fy0822/
99056953-d.html;   http://
www.loc.gov/catdir/enhancements/
fy0822/99056953-t.html.

Gould:1999:MFL

[Gou99b] Nicholas Ian Mark Gould. On modified factorizations for largescale linearly constrained optimization. SIAM Journal on Optimization, 9(4):1041– 1063, September 1999. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/29066.

Dedicated to John E. Dennis, Jr., on his 60th birthday.

Gould:2000:IMI

[Gou00] Nicholas I. M. Gould. Iterative methods for ill-conditioned linear systems from optimization. In Nonlinear optimization and related topics (Erice, 1998), volume 36 of Appl. Optim., pages 123–141. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000.

Gould:2003:SRC

[Gou03] Nicholas I. M. Gould. Some reflections on the current state of active-set and interior-point methods for constrained optimization. SIAG/OPT Viewsand-News, 14(1):2–7, April

2003. URL http://www.mcs.
anl.gov/~leyffer/views/14-
1.pdf.

Gould:2008:HGP

[Gou08] Nicholas I. M. Gould. How good are projection methods for convex feasibility problems? Computational optimization and applications, 40(1):1–12, 2008. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Gould:2012:HGE

[Gou12] Nicholas I. M. Gould. How good are extrapolated bi-projection methods for linear feasibility problems? Computational optimization and applications, 51 (3):1089–1095, 2012. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Gould:2012:URP

[GPT12] N. I. M. Gould, M. Porcelli, and P. L. Toint. Updating the regularization parameter in the adaptive cubic regularization algorithm. Computational optimization and applications, 53(1): 1–22, 2012. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Gould:1989:NCP

[GR89] Nicholas I. M. Gould and John K. Reid. New crash procedures for large systems of linear constraints. Mathematical Programming, 45(1–3):475–501, August 1989. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:2010:SDSa

[GR10a] Nicholas I. M. Gould and Daniel P. Robinson. A second derivative SQP method: Global convergence. SIAM Journal on Optimization, 20(4):2023–2048, ???? 2010. CODEN SJOPE8.

ISSN 1052-6234 (print), 1095-7189 (electronic).

Gould:2010:SDSb

[GR10b] Nicholas I. M. Gould and Daniel P. Robinson. A second derivative SQP method: Local convergence and practical issues. SIAM Journal on Optimization, 20(4):2049–2079, ???? 2010. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Gould:2012:SDS

[GR12] Nicholas I. M. Gould and Daniel P. Robinson. A secondderivative SQP method with a ‘trust-region-free’ predictor step. IMA Journal of Numerical Analysis, 32(2):580–601, April 2012. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (elec-

tronic). URL http://imajna.
oxfordjournals.org/content/
32/2/580.full.pdf+html.

Gould:2010:STR

[GRT10] Nicholas I. M. Gould, Daniel P.

Robinson, and H. Sue Thorne. On solving trust-region and other regularised subproblems in optimization. Mathematical Programming Computation, 2(1): 21–57, 2010. CODEN ???? ISSN 1867-2949 (print), 1867-2957 (electronic). URL http://www.

springerlink.com/openurl.
asp?genre=journal%26issn=
1867-295.

Gould:1998:SAI

[GS98] Nicholas I. M. Gould and Jennifer A. Scott. Sparse approximate-inverse preconditioners using norm-minimization techniques. SIAM Journal on Scientific Computing, 19(2):605– 625, 1998. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/28842.

Gould:2003:CRN

[GS03] N. I. M. Gould and J. A. Scott. Complete results from a numerical evaluation of HSL packages for the direct-solution of large sparse, symmetric linear systems of equations. Numerical Analysis Group Internal Report 2003-2, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2003.

Gould:2004:NEH

[GS04] Nicholas I. M. Gould and Jennifer A. Scott. A numerical evaluation of HSL packages for the direct solution of large sparse, symmetric linear systems of equations. ACM Transactions on Mathematical Software, 30(3): 300–325, September 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2009:SAS

[GS09] N. I. M. Gould and V. Simoncini. Spectral analysis of saddle point matrices with indefinite leading blocks. SIAM Journal on Matrix Analysis and Applications, 31(3): 1152–1171, 2009. CODEN SJ-MAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Gould:2010:SAS

[GS10] N. I. M. Gould and V. Simoncini.

Spectral analysis of saddle point matrices with indefinite leading blocks. SIAM Journal on Matrix Analysis and Applications, 31(3): 1152–1171, ???? 2010. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Gould:2007:NES

[GSH07] Nicholas I. M. Gould, Jennifer A.

Scott, and Yifan Hu. A numerical evaluation of sparse direct solvers for the solution of large sparse symmetric linear systems of equations. ACM Transactions on Mathematical Software, 33(2): 1–32, June 2007. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2005:FTR

[GST05] Nick I. M. Gould, Caroline Sainvitu, and Philippe L. Toint. A filter-trust-region method for unconstrained optimization. SIAM Journal on Optimization, 16(2): 341–357, ???? 2005. CO-DEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/60385.

Gould:1999:NCB

[GT99] Nicholas I. M. Gould and Philippe L. Toint. A note on the convergence of barrier algorithms to second-order necessary points. Mathematical Programming, 85(2, Ser. A):433– 438, 1999. CODEN MHPGA4.

ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:2000:SML

[GT00] Nicholas I. M. Gould and Philippe L. Toint. SQP methods for large-scale nonlinear programming. In System modelling and optimization (Cambridge, 1999), pages 149–178. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000.

Gould:2002:IWS

[GT02a] Nicholas I. M. Gould and Philippe L. Toint. An iterative working-set method for largescale nonconvex quadratic programming. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):109–128, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). 19th Dundee Biennial Conference on Numerical Analysis (2001).

Gould:2002:NML

[GT02b] Nicholas I. M. Gould and Philippe L. Toint. Numerical methods for large-scale nonconvex quadratic programming. In Trends in industrial and applied mathematics (Amritsar, 2001), volume 72 of Appl. Optim., pages 149–179. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002.

Gould:2003:FFF

[GT03a] N. I. M. Gould and Ph. L. Toint.

FILTRANE, a Fortran 95 filtertrust-region package for solving nonlinear feasibility problems. Technical Report RAL-TR-2003-017, Rutherford Appleton Laboratory, Chilton, Oxon, England, 2003.

Gould:2003:FIA

[GT03b] N. I. M. Gould and Philippe L.

Toint. The filter idea and its application to the nonlinear feasibility problem. In David F. Griffiths and G. Alistair Watson, editors, Proceedings of the 20th Biennial Conference on Numerical Analysis, 24–27 June, 2003, University of Dundee, Scotland, volume NA/217 of Numerical analysis report, pages 73–79. ????, ????, 2003. LCCN QA297 .D85

2003. URL http://www.maths.
dundee.ac.uk/~naconf/.

Gould:2003:GCH

[GT03c] Nick Gould and Philippe L.

Toint. Global convergence of a hybrid trust-region SQP-filter algorithm for general nonlinear programming. In System modeling and optimization, XX (Trier, 2001), volume 130 of IFIP Int. Fed. Inf. Process., pages 23– 54. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003.

Gould:2004:HMN

[GT04a] Nick Gould and Philippe L.

Toint. How mature is nonlinear optimization? In Applied mathematics entering the 21st century, pages 141–161. SIAM Press, Philadelphia, PA, USA, 2004.

Gould:2004:PQP

[GT04b] Nick Gould and Philippe L. Toint. Preprocessing for quadratic programming. Mathematical Programming, 100(1, Ser. B):95–132, 2004. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:2006:GCN

[GT06] Nicholas I. M. Gould and Philippe L. Toint. Global convergence of a non-monotone trustregion filter algorithm for nonlinear programming. In Multiscale optimization methods and applications, volume 82 of Nonconvex Optim. Appl., pages 125–150. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006.

Gould:2007:FFF

[GT07] Nicholas I. M. Gould and Philippe L. Toint. FILTRANE, a Fortran 95 filter-trust-region package for solving nonlinear least-squares and nonlinear feasibility problems. ACM Transactions on Mathematical Software, 33(1):3:1–3:23, March 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gould:2010:NPP

[GT10] N. I. M. Gould and Ph. L. Toint. Nonlinear programming without a penalty function or a filter. Mathematical Programming, 122 (1, Ser. A):155–196, 2010. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Gould:2012:ENP

[GT12] N. I. M. Gould and Ph. L. Toint.

Erratum to: Nonlinear programming without a penalty function or a filter [mr2533756]. Mathematical Programming, 131(1-2, Ser. A):403–404, 2012. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

Keller:2000:CPI

[KGW00] Carsten Keller, Nicholas I. M.

Gould, and Andrew J. Wathen. Constraint preconditioning for indefinite linear systems. SIAM Journal on Matrix Analysis and Applications, 21(4):1300–1317, 2000. CO-DEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/35180.

Scott:2004:ESD

[SHG04] Jennifer A. Scott, Yifan Hu, and Nicholas I. M. Gould. An evaluation of sparse direct symmetric solvers: an introduction and preliminary findings. In Dongarra et al. [DMW06], pages 818–827. CODEN LNCSD9. ISBN 3-540-29067-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.
springerlink.com/content/
978-3-540-29067-4;   http:/

                                        /www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-
                                        9743&volume=3732.